OpenStudy (anonymous):

a-4 / 3(a-1) - 3 / 8 = 0 (fractions) a=??

OpenStudy (anonymous):

\[\frac{23}{8} \]

OpenStudy (anonymous):

From WolframAlpha.com\[a-\frac{4}{3} (a-1)-\frac{3}{8}=0 \]\[a\text{ = }\frac{23}{8}\text{ $\approx $ }2.875 \]

OpenStudy (anonymous):

apparently -23 was the answer. thnx for helping though!!

OpenStudy (anonymous):

(a-4 )/( 3(a-1)) - 3 / 8 = 0 this is the form that gives the answer -23

OpenStudy (anonymous):

The expression began with a - 4/3 . The division has to be evaluated before the subtraction by general agreement, precedent rules. So a-(4/3), not (a-4)/3 . If the problem author intended otherwise, she should have written (a-4)/3.